A garden measuring 12 meters by 16 meters is to have a pedestrian pathway installed all around it, increasing the total area to 285 square meters. What will be the width of the pathway?

I know that the area of the path is 93, but what do I do with that???

Answer by stanbon(75887)

A garden measuring 12 meters by 16 meters is to have a pedestrian pathway installed all around it, increasing the total area to 285 square meters. What will be the width of the pathway?
--------
Draw a rectangle within a rectangle.
The dimensions of the inner rectangle are 12 by 16 meters
The dimensions of the outer rectangle are 12+2x by 16+2x
--------
Equations:
(12+2x)(16+2x) = 285 sq. meters
-----
4(6+x)(8+x) = 285
4[x^2 + 14x + 48) = 285
-----
x^2 + 14x + 48 = 71.25
----
x^2 + 14x - 23.25 = 0
-----
x = 1.5 ft is the width of the pathway.
--------------
Cheers,
Stan H.
==============